OpenStudy (anonymous):

A cone has a volume of 6 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? (1 point) Select one: a. 12 in3 b. 18 in3 c. 24 in3 d. 36 in3

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@ganeshie8 @nincompoop

OpenStudy (anonymous):

volume of cone v=1/3 pi r^2h volume of cylinder= pi r^2h=3*1/3 pi r^2h=3v=3*6=?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

yes